OpenStudy (anonymous):

Bag A has 4 blue and 5 red balls. Bag B has 5 blue and 4 red balls. I randomly choose a ball from bag A, then putting it into B. I then randomly choose a ball from bag B. What is the probability that the ball chosen from bag B is red?

OpenStudy (anonymous):

is it (4/9*4/10 +5/9* 5/10 = 41/90)=41/90

OpenStudy (kropot72):

Probability of having a fifth red ball in bag B is 5/9. Probability that the ball chosen from bag B is red is: \[\frac{1}{10}(4+\frac{5}{9})\]=41/90